Solve for L: S=2WL+2WH+2LH
Can you please explain the steps in the easiest way possible? I'm in Algebra and i have a HUGE test coming up so feel free to leave other different methods to solve the problem fast and other sites where i can get th help i need. Thanks

Answer by Alan3354(69443)

Solve for L: S=2WL+2WH+2LH
Can you please explain the steps in the easiest way possible? I'm in Algebra and i have a HUGE test coming up so feel free to leave other different methods to solve the problem fast and other sites where i can get th help i need. Thanks
----------------
"Solve for" means to get it one side of the equal sign by itself.
S=2WL+2WH+2LH
S-2WH = 2WL+2LH
S-2WH = L*(2W+2H)
L = (S-2WH)/(2W+2H)